OpenStudy (anonymous):

MEDAL AND FAN How many solutions does the equation have? a + 5 = 1/5(5a +25) A. One Solution B. Infinite Solutions C. No Solution

OpenStudy (anonymous):

a+5=1/5(5a+25) a+5=1a+5

OpenStudy (anonymous):

so its b

OpenStudy (anonymous):

wait so how did u know tht

OpenStudy (anonymous):

@cody7744 so how did u know tht

OpenStudy (anonymous):

hey, just so you know, you're in the wrong section for this question. it should be in a math section, and this is english. i think you'd be able to find some better help in the mathematics subject than here.

OpenStudy (anonymous):